OpenStudy (anonymous):

essica stands 150 feet from the base of a tall building. she measures the angle the angle from her eye to the top of the building to be 84 degrees. If her eye level is five feet above the ground, how tall is the building ?

OpenStudy (anonymous):

w = 150 theta = 84 distance from ground is 5.. lets keep it in mind h = x + 5 (where x is height of peices she's looking at) tan84 = (x)/150 x=1427.15 ft h = 1427.15 + 5 = 1432.15 ft